OpenStudy (anonymous):

I am just not good at subtracting radicals.... sqrt{5}-sqrt[3]{20}-sqrt[3]{5}-sqrt[2]{20}

OpenStudy (anonymous):

\[\sqrt{5}-\sqrt[3]{20}-\sqrt[3]{5}-\sqrt[2]{20} \]

OpenStudy (anonymous):

Thanks again Kboy!

OpenStudy (anonymous):

\[\sqrt{20}=2\sqrt{5}\]...are u sure its \[\sqrt[3]{20}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can rewrite as (sqrt(5) - 2sqrt(5)) to get -sqrt(5), but i cant see much that u cn do with the other two terms.

OpenStudy (anonymous):

Yes and I checked and that is exactly how the question is written

OpenStudy (anonymous):

ugh lol!

OpenStudy (anonymous):

ok.lets do this:\[\sqrt[3]{20}=\sqrt[3]{5*4}=\sqrt[3]{5}*\sqrt[3]{4}\]

OpenStudy (anonymous):

now take out \[\sqrt[3]{5}\] as a like term

OpenStudy (anonymous):

so the answer is \[-\sqrt{5} -\sqrt[3]{5}(1 + \sqrt[3]{4})\]

OpenStudy (anonymous):

ok... so do you eliminate \[\sqrt[3]{5}\] from the equation

OpenStudy (anonymous):

not eliminate...extract as a common term

OpenStudy (anonymous):

ok so I am trying to decifer this can you please go step by step with me sorry for asking this of you

OpenStudy (anonymous):

ok first rearrange the initial equation by grouping "like terms".we have square roots and cube roots. so we rewrite initial equation as:\[(\sqrt{5}-\sqrt{20})-(\sqrt[3]{20}+\sqrt[3]{5})\]

OpenStudy (anonymous):

ok but you are suppose to subtract everything?

OpenStudy (anonymous):

yes,we are still. now we rewrite \[\sqrt{20} \] as \[2\sqrt{5}\] and \[\sqrt[3]{20} \] as \[\sqrt[3]{4}*\sqrt[3]{5}\]

OpenStudy (anonymous):

ok but you are suppose to subtract everything?

OpenStudy (anonymous):

why did you add \[\sqrt[3]{5}\] then?

OpenStudy (anonymous):

we then get:\[(\sqrt{5}-2\sqrt{5}) - (\sqrt[3]{5} + \sqrt[3]{5}*\sqrt[3]{4})\]

OpenStudy (anonymous):

the negative sign outside the bracket will make it negative if you multiply out, so we are still subtracting it

OpenStudy (anonymous):

ok but you are suppose to subtract everything?

OpenStudy (anonymous):

ohhhhh ok

OpenStudy (anonymous):

now we simplify and get:\[(-\sqrt{5})-(\sqrt[3]{5}+\sqrt[3]{5}*\sqrt[3]{4})\]

OpenStudy (anonymous):

now do you see that you can take out \[\sqrt[3]{5}\] as a common factor in the second bracket

OpenStudy (anonymous):

but you are adding it not dividing?

OpenStudy (anonymous):

we are subtracting, the negative sighn outside the bracket will make it negative agen

OpenStudy (anonymous):

ohhh ok i see

OpenStudy (anonymous):

the answer is :\[-\sqrt{5}-\sqrt[3]{5}(1 + \sqrt[3]{4})\]

OpenStudy (anonymous):

ok you lost me I thought it would\[\sqrt{5}*\sqrt[3]{4}\]

OpenStudy (anonymous):

and couldn't you simplify simplify the square rooted 5's

OpenStudy (anonymous):

lol. mayb capslock will help...WE ARE SUBTRACTING. take the answer i gave you, it is right.and no u cant simplify the sqrt(5). if u dnt undertand im sure the lecturer shud b able to clarify things further for u.

OpenStudy (anonymous):

ok ok im sorry thank you for everything

OpenStudy (anonymous):

no worries.sorry i cudnt make you fully understand

OpenStudy (anonymous):

it's alright it's my fault

OpenStudy (anonymous):

just practice them more,there isnt much to it,im sure ul master them soon.goodluck. :)

OpenStudy (anonymous):

haha thanks again